Give an upper bound and a lower bound for the expression 1/(a^(4)+3a^(2)+1) if a∈R


1
Expert's answer
2021-02-04T06:31:53-0500

Solution

Let f(x) = 1/(x4+3x2+1)

Domain of f(x) is (-∞; ∞)

f(x) is continuous, f(x) >0 for x∈ (-∞; ∞)

f(x) is symmetric about y-axis

lim|x|->∞ f(x) = 0 => axis Ox is asymptote of f(x)

df/dx = -(4x3+6x)/(x4+3x2+1) => df/dx = 0 only when x=0

d2f/dx2 = -((12x2+6) (x4+3x2+1)- (4x3+6x)2)/(x4+3x2+1)2 ; for x = 0  d2f/dx2 = -6 <0 => x = 0 is the point of maximum

f(0) = 1

So range of f(x) is 0 < y ≤ 1

Therefore lower bound is 0, upper bound is 1

Answer

lower bound is 0, upper bound is 1